Jonathan’s Enterprises purchases $4,562,500 in goods per year from its sole supplier on terms of 2/15, net 50. If the firm chooses to pay on time but does not take the discount, what is the effective annual percentage cost of its non-free trade credit? (Assume a 365-day year.)


1
Expert's answer
2021-05-31T15:56:02-0400

Jonathan's Enterprises purchases $4,562,500 goods per year 2/15, net 50 means that company will get 2% discount if it pays within 15 days, otherwise it has to pay the full amount within 50 days.


"2\%\ of\ \$4,562,500= \$91,250 \ \ discount"

Therefore company can pay = $4,562,500 - $91250 = $4471250 in 15 days

Or $4,562,500 in 50 days


The difference is 35 days, therefore we need to calculate the interest rate over the 35 days and then calculate the Effective Annual Rate associated with that 35 days interest rate.


35 days interest rate "r = \dfrac{\$91,250}{ \$4,471,250} = 0.02041,\ or\ 2.041\%"

Therefore for a year or 365 days periods n = 365/35 = 10.43


The effective annual rate

"EAR= (1+r) ^n -1"

"= (1+ 2.041) ^{10.43} -1"

"= 0.2346\ or \ 23.46%"


Therefore, the effective annual percentage cost of its non-free trade credit is 23.46%